# LightStereo: Channel Boost Is All Your Need for Efficient 2D Cost Aggregation

### v1

@@ 4.5 Runtime Analysis
As shown in Table 6, it is evident that the LightStereo framework demonstrates commendable efficiency across its constituent modules. Each component, including feature extraction, cost computation, cost aggregation, and disparity regression, contributes to the overall computational time. It is evident that as the number of blocks increases, the time required for cost aggregation in LightStereo-S, LightStereo-M, and LightStereo-L also increases. For LightStereo-L*, the time spent on feature extraction is extended due to the adoption of a more complex EfficientNetV2. These runtime breakdowns underscore the effectiveness of the LightStereo framework in achieving real-time stereo-matching capabilities.
5 Conclusion
@@
In conclusion, this paper designs a lightweight 2D encoder-decoder aggregation net to achieve precise and fast disparity estimation, called LightStereo. While similar approaches have been explored previously, our novel contribution lies in optimizing performance through a targeted emphasis on the disparity channel dimension within the 3D cost volume, which encapsulates the distribution of matching costs. Our exhaustive exploration has led to the development of numerous strategies to enhance the capacity of this crucial dimension, ensuring both precision and efficiency in disparity estimation. LightStereo offers a compelling solution for accelerating the matching process while maintaining high levels of accuracy and efficiency.
@@
Acknowledgements. This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Open Projects Program of the State Key Laboratory of Multimodal Artificial Intelligence Systems.

### v3

@@ IV-D Ablation Study
Runtime Analysis. In, Table VI, LightStereo demonstrates commendable efficiency across its constituent modules. As the number of blocks increases, the time required for cost aggregation in LightStereo-S, LightStereo-M, and LightStereo-L also increases. For LightStereo-H, the time spent on feature extraction is extended due to the adoption of a more complex EfficientNetV2. These runtime breakdowns underscore the effectiveness of the LightStereo framework in achieving real-time stereo-matching capabilities.
IV-E Comparison with State-of-the-art on Real Datasets
KITTI benchmarks. We evaluate the results of our LightStereo variants on the KITTI 2012 and 2015 online benchmarks. As illustrated in Table V, LightStereo-S runs the fastest. Notably, LightStereo-H surpasses all other lightweight state-of-the-art stereo-matching networks across every evaluation metric on KITTI 2012. Additionally, LightStereo-H achieved the best results in terms of D1-bg and D1-all among lightweight models on KITTI 2015.
Cross-domain generalization. We evaluated the generalization performance of our model on real-world datasets such as KITTI12 [26], KITTI15 [27], Middlebury [47], as shown in Table VII. LightStereo shows superior generalization over other lightweight methods, which indicates that our approach has moderate complexity to avoid overfitting.
V Conclusion
@@
This paper designs a lightweight 2D encoder-decoder aggregation network to achieve precise and fast disparity estimation, called LightStereo. While similar approaches have been explored previously, our novel contribution lies in optimizing performance through a targeted emphasis on the disparity channel dimension within the 3D cost volume, which encapsulates the distribution of matching costs. Our exhaustive exploration has led to the development of numerous strategies to enhance the capacity of this crucial dimension, ensuring both precision and efficiency in disparity estimation. LightStereo offers a compelling solution for accelerating the matching process while maintaining high levels of accuracy and efficiency.
@@
Acknowledgements. This work was supported by the National Natural Science Foundation of China under Grant 62373356 and the Joint Funds of the National Natural Science Foundation of China under U24B20162.